### v2

# Symmetry transformation group arising from the Laplace–Runge–Lenz vector

Abstract.
The Kepler problem in classical mechanics exhibits a rich structure of conserved quantities, highlighted by the Laplace–Runge–Lenz (LRL) vector. Through Noether’s theorem in reverse, the LRL vector gives rise to a corresponding infinitesimal dynamical symmetry on the kinematical variables, which is well known in the literature. However, the physically relevant part of the LRL vector is its direction angle in the plane of motion (since its magnitude is just a function of energy and angular momentum). The present work derives the infinitesimal dynamical symmetry corresponding to the direction part of the LRL vector and obtains the explicit form of the symmetry transformations that it generates. When combined with the rotation symmetries, the resulting symmetry group is shown to be the semi-direct product of $SO(3)$ and $\mathbb{R}^{3}$. This stands in contrast to the $SO(4)$ symmetry group generated by the LRL symmetries and the rotations. As a by-product, the action of the new infinitesimal symmetries on all of the conserved quantities is obtained. The results are given in terms of the physical kinematical variables in the Kepler problem, rather than in an enlarged auxiliary space in which the LRL symmetries are usually stated.
1. Introduction
The Kepler problem in Newtonian mechanics has been widely studied and encompasses both a particle moving under inverse-square force and two bodies moving under mutual gravitational attraction. One of its most interesting features is the existence of the conserved Laplace–Runge–Lenz (LRL) vector (see e.g. [1, 2] and references therein). It lies in the plane of motion, which is orthogonal to the angular momentum vector, and points in the direction given by the periapsis in the case of inverse-square motion and the position of closest separation in the case of two-body motion. This conserved vector has a long history, going back over three centuries, and plays a key role in understanding the shape and orientation of astronomical orbits in the solar system. Its quantum mechanical analog was important historically in the initial derivation of energy levels in the hydrogen atom (see e.g. [3] and references therein).
In contrast to the conservation of energy and angular momentum in the Kepler problem, which come directly from Noether’s theorem using the manifest time-translation and rotation symmetries of Newtonian space, the LRL vector does not arise from any kinematic symmetry principle. Nevertheless, Noether’s theorem in reverse [4, 5] states that every conserved quantity is generated by some infinitesimal symmetry transformation that leaves invariant the action principle for the Kepler problem. Since the LRL vector comprises three independent conserved quantities, its symmetry generator consists of three independent infinitesimal transformations, which were first worked out in Ref.[6]. They turn out to involve the velocity variable, unlike the infinitesimal time-translation and rotation transformations which involve only the time and position variables. In particular, the LRL infinitesimal symmetry transformations are primary examples of what are called dynamical symmetries, in contradistinction to point symmetries which are typified by infinitesimal time-translations and rotations. The only additional point symmetry admitted by the Kepler problem consists of a scaling of the time and position variables, but this scaling transformation does not leave the action principle invariant.
The five point symmetries together with the three LRL symmetries in the Kepler problem form an 8-dimensional Lie algebra in which the time-translation and scaling symmetries constitute a two-dimensional abelian subalgebra, namely they commute with the rotations and LRL symmetries. The latter symmetries are known to form a 6-dimensional Lie subalgebra when the LRL symmetries are normalized to have the same physical dimension as the rotations. This subalgebra’s structure turns out to be $\mathfrak{so}(4)$ when the energy is negative, and $\mathfrak{so}(3,1)$ when the energy is positive [7]. These Lie algebras are isomorphic to the symmetry algebra of isometries of, respectively, a 3-sphere $S^{3}$ in $\mathbb{R}^{4}$ and a 3-hyperboloid $H^{3}$ in Minkowski space $\mathbb{R}^{3+1}$.
One explanation for the appearance of four-dimensional spaces is provided by Moser’s observation [8] that, in the negative energy case, the Kepler problem is equivalent to non-affine geodesic motion on $S^{3}$ in a four-dimensional auxiliary space under stereographic projection. A kinematical description of the projection is given by [9, 7] identifying the stereographic coordinates as the components of the Kepler velocity variable. The rotations in the auxiliary space correspond in the Kepler problem to changing the angular momentum while keeping the energy unchanged. These rotations constitute a $SO(4)$ transformation group that combines the ordinary physical $SO(3)$ rotations and the LRL symmetry transformations. A similar equivalence [3] holds in the positive energy case by considering non-affine geodesic motion on $H^{3}$.
One remaining question is how does the group generated by the LRL infinitesimal symmetry transformations act just on the time, position, and velocity variables in the Kepler problem. It is hard to find an explicit description of this group in the literature (although a description in terms of the coordinates in the auxiliary 4-dimensional space is known [10]).
The first main result of the present work will be to derive the explicit form of the LRL symmetry transformation group in terms of the kinematic physical variables.
Another remaining question, which is not commonly discussed in the literature, arises from the magnitude of LRL vector. This quantity is given by a scalar expression in terms of the energy and magnitude of the angular momentum. The corresponding infinitesimal symmetry transformation, coming from Noether’s theorem in reverse, is just a linear combination of the generators of time-translation and rotation in the plane of motion. Namely it is a point symmetry, involving only the time and position variables. On the one hand, this choice of normalization of the LRL vector is essential for the subalgebra generated by the LRL symmetries and rotations to be $\mathfrak{so}(4)$ when energy is negative and $\mathfrak{so}(3,1)$ when energy is positive. Yet on the other hand, it is a free choice, which raises the question of whether a different or perhaps simpler structure is possible for the resulting symmetry subalgebra if some other choice of normalization is made. A natural choice is to normalize it to be a unit vector, retaining only the physically relevant direction information contained in the LRL vector. This will be called the LRL direction vector.
As a second main result, the components of the LRL direction vector will be shown to generate three infinitesimal symmetry transformations that commute with themselves. This constitutes a symmetry subalgebra $\mathbb{R}^{3}$ which is abelian and belongs to the larger 6-dimensional symmetry algebra $\mathfrak{so}(3)\rtimes\mathbb{R}^{3}$ given by a semi-direct sum with rotations $\mathfrak{so}(3)$. The explicit form of the corresponding symmetry transformation group, which is $SO(3)\rtimes\mathbb{R}^{3}$ acting on the time, position, and velocity variables in the Kepler problem, will be presented.
The rest of the paper is organized as follows. In section 2, the generators of the infinitesimal symmetry transformations given by Noether’s theorem in reverse applied to the constants of motion in the Kepler problem will be reviewed using a modern formulation. The symmetry algebras respectively containing the LRL vector and the LRL direction vector will be summarized in section 3 through the corresponding Poisson brackets by familiar use of a Hamiltonian formulation. In addition, the action of the infinitesimal symmetries on the constants of motion will be derived in terms of the Poisson brackets. In section 4, the explicit symmetry transformation group arising from the LRL direction vector will be worked out on the kinematical variables. A key part of the derivation involves the use of the infinitesimal symmetry action on the energy, angular momentum, and the LRL direction vector. In section 5, the derivation is extended to obtain the explicit form of the LRL symmetry transformation group on the kinematic variables. This turns out to be more complicated because the structure of the Poisson brackets involving the LRL vector is richer in comparison to the Poisson brackets which involve the LRL direction vector. Some concluding remarks will be given in section 6.
2. Infinitesimal symmetry generators
The equations of motion of the Kepler problem are given by
$\ddot{{\vec{r}}}=-\kappa|{\vec{r}}|^{-2}{\hat{r}}$ (2.1)
where $\kappa$ is a positive constant. For particle motion in Newtonian gravity, $\vec{r}$} is the position of the particle with respect to the central mass, while $\kappa=GM$ where $M$ is the central mass and $G$ is the gravitational constant. For two body motion under Newtonian gravitational attraction, $\vec{r}$} is the relative position of the bodies with respect to their center of mass, and $\kappa=G\mu$ where $\mu$ is the reduced mass of the two bodies. In both cases, $\hat{r}=|{\vec{r}}|^{-1}{\vec{r}}$ is the unit position vector.
The constants of motion consist of energy
$E=\tfrac{1}{2}|\dot{{\vec{r}}}|^{2}-\kappa|{\vec{r}}|^{-1},$ (2.2)
angular momentum
$\vec{L}={\vec{r}}\times\dot{{\vec{r}}},$ (2.3)
and the LRL vector
$\vec{A}=\dot{{\vec{r}}}\times\vec{L}-\kappa{\hat{r}}$. (2.4)
Here these quantities have been normalized by dividing out the particle mass in the case of inverse-square motion and the reduced mass in the case of two-body motion.
The LRL vector is orthogonal to the angular momentum, $\vec{A}\cdot\vec{L}=0$, and has magnitude $|\vec{A}|=\sqrt{\kappa^{2}+2E|\vec{L}|^{2}}$. For circular orbits, $\vec{A}$ vanishes, whereas for elliptical, parabolic, and hyperbolic orbits, $\vec{A}$ is non-trivial and points in the direction of the periapsis. Its magnitude is proportional to the eccentricity of the orbit, $e=|\vec{A}|/\kappa$. Therefore, among the three components of $\vec{A}$, only the angle of $\vec{A}$ in the plane of motion for non-circular orbits represents an independent constant of motion. If this angle, $\Theta$, is measured relative to an axis defined by the periapsis point in the orbit, then it is given by $\Theta=0$.
Thus, $E$, $\vec{L}$, $\Theta$ provide 5 independent constants of motion. This is the maximal number for any type of motion in three dimensions. There is an additional conserved quantity, which involves $t$ and is not constant of motion, that can be defined as the time $T$ at which the periapsis point is reached in a non-circular orbit. An explicit formula for $T$, along with discussion of its physical significance, is given in Ref.[11] for general central force dynamics.
Noether’s theorem gives a one-to-one correspondence between conserved quantities and infinitesimal dynamical symmetries of an action principle. For the Kepler problem (2.1), the action principle is $S=\int_{t_{1}}^{t_{2}}\mathcal{L}({\vec{r}},\dot{{\vec{r}}})\,dt$ where
$\mathcal{L}({\vec{r}},\dot{{\vec{r}}})=\tfrac{1}{2}|\dot{{\vec{r}}}|^{2}+\kappa|{\vec{r}}|^{-1}$ (2.5)
is the Lagrangian, whose variational derivative yields the equations of motion,
$\frac{\partial{\mathcal{L}}}{\partial{{\vec{r}}}}-\frac{d}{dt}\frac{\partial{\mathcal{L}}}{\partial{\dot{{\vec{r}}}}}=-\ddot{{\vec{r}}}-\kappa|{\vec{r}}|^{-2}{\hat{r}}$. (2.6)
Recall that adding a total time derivative to $\mathcal{L}$, which just changes $S$ by the addition of an endpoint term, does not affect the equations of motion. In this context, an infinitesimal dynamical symmetry is a generator
${\bf X}={\vec{P}}(t,{\vec{r}},\dot{{\vec{r}}})\cdot\partial_{{\vec{r}}},$ (2.7)
acting on the dynamical variable $\vec{r}$}, such that the action principle is invariant up to an endpoint term. Specifically, the condition of invariance is given by ${\rm pr}{\bf X}(S)=W|_{t_{1}}^{t_{2}}$ for some function $W(t,{\vec{r}},\dot{{\vec{r}}})$, where
${\rm pr}{\bf X}={\vec{P}}\cdot\partial_{{\vec{r}}}+\dot{{\vec{P}}}\cdot\partial_{\dot{{\vec{r}}}}$ (2.8)
denotes the prolongation of $\bf X$}, acting on the Lagrangian variables $\vec{r}$} and $\dot{{\vec{r}}}$. Infinitesimal symmetry invariance can be usefully restated entirely in terms of the Lagrangian as
${\rm pr}{\bf X}(\mathcal{L})=\frac{\partial{\mathcal{L}}}{\partial{{\vec{r}}}}\cdot{\vec{P}}+\frac{\partial{\mathcal{L}}}{\partial{\dot{{\vec{r}}}}}\cdot\dot{{\vec{P}}}=\frac{d{W}}{d{t}}.$ (2.9)
Proposition 1.
A function $C(t,{\vec{r}},\dot{{\vec{r}}})$ is a conserved quantity for the Kepler problem, namely $\dot{C}=0$ for all solutions of the equations of motion (2.1), if and only if the generator ${\bf X}={\vec{P}}(t,{\vec{r}},\dot{{\vec{r}}})\cdot\partial_{{\vec{r}}}$ is an infinitesimal dynamical symmetry of the Lagrangian (2.5), where
${\vec{P}}=\dfrac{\partial C}{\partial\dot{{\vec{r}}}}.$ (2.10)
Any infinitesimal dynamical symmetry gives rise to a (one-parameter) transformation group of dynamical symmetries which simply represents the flow generated by the vector field ${\rm pr}{\bf X}$. This can be understood in two different ways.
The first way is to work in the space of Lagrangian variables $({\vec{r}},\dot{{\vec{r}}})$, or equivalently, in phase space (Hamiltonian) variables $({\vec{r}},{\vec{v}})$, where ${\vec{v}}=\dot{{\vec{r}}}$ is the velocity vector. Let $\mathcal{D}_{t}$ denote time derivatives evaluated through use of the equations of motion (2.1), in particular $\mathcal{D}_{t}{\vec{r}}={\vec{v}}$ and $\mathcal{D}_{t}{\vec{v}}=-\kappa|{\vec{r}}|^{-2}{\hat{r}}$, whereby it acts as an operator on phase space,
$\mathcal{D}_{t}={\vec{v}}\cdot\partial_{{\vec{r}}}-\kappa|{\vec{r}}|^{-2}{\hat{r}}\cdot\partial_{{\vec{v}}}.$ (2.11)
Then the resulting symmetry transformation group $({\vec{r}},{\vec{v}})\to({\vec{r}},{\vec{v}})^{*}$ is given by
$({\vec{r}},{\vec{v}})^{*}=\exp(\epsilon\,\hat{\bf X})({\vec{r}},{\vec{v}}),\quad\hat{\bf X}={\vec{P}}\cdot\partial_{{\vec{r}}}+\mathcal{D}_{t}{\vec{P}}\cdot\partial_{{\vec{v}}}$ (2.12)
with $\varepsilon$ being the group parameter, where $\varepsilon=0$ yields the identity transformation. Note that this transformation group keeps $t$ fixed.
A second way to define an equivalent transformation group is by having the transformations act directly on solutions $({\vec{r}}(t),{\vec{v}}(t))$ of the equations of motion (2.1). This yields
$({\vec{r}}(t),{\vec{v}}(t))^{*}=\exp(\varepsilon\,{\rm pr}{\bf X})({\vec{r}}(t),{\vec{v}}(t))$ (2.13)
where, now,
${\rm pr}{\bf X}={\vec{P}}(t,{\vec{r}}(t),{\vec{v}}(t))\cdot\partial_{{\vec{r}}}+{\vec{P}}(t,{\vec{r}}(t),{\vec{v}}(t))\dot{\mathstrut}\cdot\partial_{{\vec{v}}}$ (2.14)
using the ordinary time derivative restricted to solutions.
Both transformation groups (2.12) and (2.13) represent the one-parameter dynamical symmetry group generated by $\bf X$}.
A special case of dynamical symmetry groups comprises point symmetry groups (see e.g. Ref.[12, 4]) which are commonly expressed by point transformations
${\vec{r}}\to{\vec{r}}{}^{*}={\vec{r}}+\varepsilon\vec{\xi}+O(\varepsilon^{2}),\quad{\vec{v}}\to{\vec{v}}{}^{*}={\vec{v}}+\varepsilon\dot{\vec{\xi}}+O(\varepsilon^{2}),\quad t\to t{}^{*}=t+\varepsilon\tau+O(\varepsilon^{2})$ (2.15)
where $\vec{\xi}$ and $\tau$ are functions only of $t$ and $\vec{r}$}. Their infinitesimal generator is given by
${\bf Y}_{\text{point}}=\tau\partial_{t}+\vec{\xi}\cdot\partial_{{\vec{r}}}+\dot{\vec{\xi}}\cdot\partial_{{\vec{v}}},$ (2.16)
corresponding to the transformation group $(t,{\vec{r}},{\vec{v}})^{*}=\exp(\varepsilon\,{\bf Y}_{\text{point}})(t,{\vec{r}},{\vec{v}})$. These symmetries act in the coordinate space $(t,{\vec{r}},{\vec{v}})$, unlike dynamical symmetries which act in phase space $({\vec{r}},{\vec{v}})$ or on solutions of the equations of motion.
Nevertheless, point symmetries naturally give rise to a transformation group acting in phase space. The transformations are readily derived by considering the effect of a point symmetry on solutions of the equations of motion (2.1), where each solution geometrically describes a curve $({\vec{r}}(t),{\vec{v}}(t))$ in the coordinate space $(t,{\vec{r}},{\vec{v}})$. Under a point symmetry, the curve is mapped to another curve $({\vec{r}}^{*}(t^{*}),{\vec{v}}^{*}(t^{*}))$ in the same space. This mapping has an equivalent formulation in which $t$ is kept fixed and only $({\vec{r}}(t),{\vec{v}}(t))$ is changed: the resulting infinitesimal transformation has the form [12]
${\vec{r}}{}^{*}(t)={\vec{r}}(t)+\varepsilon\big(\vec{\xi}(t,{\vec{r}}(t))-\tau(t,{\vec{r}}(t))\,{\vec{v}}(t)\big)+O(\varepsilon^{2}),$ (2.17)
${\vec{v}}{}^{*}(t)={\vec{v}}(t)+\varepsilon\big(\vec{\xi}(t,{\vec{r}}(t))-\tau(t,{\vec{r}}(t))\,{\vec{v}}(t)\big)\dot{\mathstrut}+O(\varepsilon^{2}).$
Its corresponding infinitesimal symmetry generator has the following simple formulation in phase space,
$\hat{\bf X}_{\text{point}}={\vec{P}}_{\text{point}}\cdot\partial_{{\vec{r}}}+\mathcal{D}_{t}{\vec{P}}_{\text{point}}\cdot\partial_{{\vec{v}}},\quad{\vec{P}}_{\text{point}}=\vec{\xi}-\tau{\vec{v}}.$ (2.18)
Observe that this generator (2.18) is related to the previous generator (2.16) by ${\bf Y}_{\text{point}}=\hat{\bf X}_{\text{point}}+\tau D_{t}$, where
$D_{t}=\partial_{t}+{\vec{v}}\cdot\partial_{{\vec{r}}}+\mathcal{D}_{t}{\vec{v}}\cdot\partial_{{\vec{v}}}.$ (2.19)
In particular, when acting on solutions $({\vec{r}}(t),{\vec{v}}(t))$, the vector field $\tau D_{t}$ represents a trivial infinitesimal symmetry that leaves every solution unchanged.
Consequently, what distinguishes a point symmetry generator from a dynamical symmetry generator is their dependence on $\vec{v}$}.
Proposition 2.
A point symmetry generator ${\vec{P}}_{\text{point}}(t,{\vec{r}},{\vec{v}})$ is strictly linear in $\vec{v}$}, namely,
$\dfrac{\partial{\vec{P}}_{\text{point}}}{\partial{\vec{v}}}=\tau(t,{\vec{r}})\mathbf{I}$ (2.20)
where $\mathbf{I}$ denotes the identity matrix. A genuine dynamical symmetry generator ${\vec{P}}_{\text{dyn.}}(t,{\vec{r}},{\vec{v}})$ is either non-strictly linear or nonlinear in $\vec{v}$}, namely,
$\dfrac{\partial{\vec{P}}_{\text{dyn.}}}{\partial{\vec{v}}}=\mathbf{f}(t,{\vec{r}},{\vec{v}})$ (2.21)
with the matrix function $\mathbf{f}(t,{\vec{r}},{\vec{v}})$ satisfying
$\frac{\partial{\mathbf{f}}}{\partial{{\vec{v}}}}=0,\quad\mathbf{f}\neq\tau\mathbf{I}$ (2.22a)
or
$\frac{\partial{\mathbf{f}}}{\partial{{\vec{v}}}}\neq 0.$ (2.22b)
An important observation is that any infinitesimal dynamical symmetry (2.7) can be extended to act in the coordinate space $(t,{\vec{r}},{\vec{v}})$ by means of the generator
${\bf Y}=\hat{\bf X}+\tau D_{t}=\tau\partial_{t}+\big({\vec{P}}+\tau{\vec{v}}\big)\cdot\partial_{{\vec{r}}}+\big(\mathcal{D}_{t}{\vec{P}}-\tau|{\vec{r}}|^{-2}{\hat{r}}\big)\cdot\partial_{{\vec{v}}}$ (2.23)
where $\tau$ is a completely arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. The flow of this vector field (2.23) generates the transformation group
$(t,{\vec{r}},{\vec{v}})\to(t,{\vec{r}},{\vec{v}})^{*}=\exp(\varepsilon\,{\bf Y})(t,{\vec{r}},{\vec{v}})$ (2.24)
with parameter $\varepsilon$. In particular,
$t{}^{*}=t+\varepsilon\tau+O(\varepsilon^{2}),\quad{\vec{r}}{}^{*}={\vec{r}}+\varepsilon({\vec{P}}+\tau{\vec{v}})+O(\varepsilon^{2}),\quad{\vec{v}}{}^{*}={\vec{v}}+\varepsilon(\mathcal{D}_{t}{\vec{P}}-\tau|{\vec{r}}|^{-2}{\hat{r}})+O(\varepsilon^{2}).$ (2.25)
In this transformation group, $\tau$ constitutes a gauge freedom, since both ${\rm pr}{\bf X}$ and $\bf Y$} act in an identical way on solutions $({\vec{r}}(t),{\vec{v}}(t))$.
Existence of this gauge freedom has not been widely recognized in the literature. However, it will turn out to play a key role in finding the explicit form of dynamical symmetry transformations generated from the LRL vector
2.1. Infinitesimal symmetry generators from the constants of motion
From Proposition 1, a one-to-one correspondence exists between constants of motion and time-independent infinitesimal dynamical symmetries for the Kepler problem. This correspondence is given explicitly by the relation (2.10). For notational ease, index notation will be adopted hereafter.
Energy (2.2) and angular momentum (2.3) respectively give
$P^{i}[E]=\frac{\partial{E}}{\partial{\dot{r}^{i}}}=\dot{r}^{i},\qquad{\bf X}_{E}=\dot{r}^{i}\partial_{r^{i}}$ (2.26)
and
$P^{i}[L^{j}]=\frac{\partial{L^{j}}}{\partial{\dot{r}^{i}}}=\epsilon_{ijk}r^{k},\qquad{\bf X}_{L^{j}}=\epsilon_{ijk}r^{k}\partial_{r^{i}}$ (2.27)
(where indices are raised and lowered by the Euclidean metric $\delta_{ij}$). Notice these expressions are strictly linear in $\dot{r}^{i}=v^{i}$, namely $\dfrac{\partial^{2}E}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=\delta_{ik}$ and $\dfrac{\partial^{2}L^{j}}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=0$ are proportional to the identity matrix. Hence, they represent generators of point symmetries.
For energy, the generator ${\bf X}_{E}$ produces the infinitesimal symmetry transformation
$r^{i}\to r^{i}{}^{*}=r^{i}+\varepsilon\dot{r}^{i}+O(\varepsilon^{2})$ (2.28)
with group parameter $\varepsilon$. The resulting group of symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ is given by
$r^{i}(t){}^{*}$ $=\exp(\varepsilon\hat{\bf X}_{E})r^{i}(t)=\exp\Big(\varepsilon\frac{d}{dt}\Big)r^{i}(t)=r^{i}(t+\epsilon),$ (2.29)
$v^{i}(t){}^{*}$ $=\exp(\varepsilon\hat{\bf X}_{E})v^{i}(t)=\exp\Big(\varepsilon\frac{d}{dt}\Big)v^{i}(t)=v^{i}(t+\epsilon)$
where $\hat{\bf X}_{E}=\dot{r}^{i}(t)\partial_{r^{i}}+\dot{v}^{i}(t)\partial_{v^{i}}$. This is clearly a time-translation symmetry group. In particular, it is equivalent to the point transformation
$t\to t^{*}=t-\varepsilon,\quad r^{i}\to r^{i}{}^{*}=r^{i},\quad v^{i}\to v^{i}{}^{*}=v^{i}$ (2.30)
acting in the coordinate space $(t,r^{i},v^{i})$. Note that the minus sign in this transformation comes from the general relation between a point transformation (2.17) and its generator (2.18) on phase space, which shows that $\tau=-1$ and $\xi^{i}=0$.
For angular momentum, the generator ${\bf X}_{L^{j}}$ produces the infinitesimal symmetry transformation
$r^{i}\to r^{i}{}^{*}=r^{i}+\epsilon^{ijk}\varepsilon_{j}r_{k}+O(\varepsilon^{2})$ (2.31)
with group parameter $\varepsilon^{j}$ (which is a vector). This is clearly an infinitesimal rotation acting on $r^{i}$. Note that its generator (2.18) has $\tau=0$ and $\xi^{i}=\epsilon^{ijk}r_{k}$. The corresponding symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ are given by the rotation group
$r^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{L^{j}})r^{i}(t)=R^{i}{}_{j}(\varepsilon)r^{j}(t),\quad v^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{L^{j}})v^{i}(t)=R^{i}{}_{j}(\varepsilon)v^{j}(t)$ (2.32)
with
$R^{i}{}_{j}(\varepsilon)=\cos|\varepsilon|\ \delta^{i}{}_{j}+(1-\cos|\varepsilon|)\hat{\varepsilon}^{i}\hat{\varepsilon}^{j}+\sin|\varepsilon|\ \hat{\varepsilon}_{k}\epsilon^{ki}{}_{j},\quad|\varepsilon|=\varepsilon^{j}\varepsilon_{j},\quad\hat{\varepsilon}^{i}=|\varepsilon|^{-1}\varepsilon^{i}$ (2.33)
being the familiar rotation operator, where $\hat{\bf X}_{L}=\epsilon_{ijk}r^{k}(t)\partial_{r^{i}}+\epsilon_{ijk}v^{k}(t)\partial_{v^{i}}$. In particular, the rotation angle is $|\varepsilon|$ and the rotation axis is $\hat{\varepsilon}^{i}$, determined by the group parameter $\varepsilon^{i}$. These symmetries acts as a point transformation in the coordinate space $(t,r^{i},v^{i})$:
$t\to t^{*}=t,\quad r^{i}\to r^{i}{}^{*}=R^{i}{}_{j}(\varepsilon)r^{j},\quad v^{i}\to v^{i}{}^{*}=R^{i}{}_{j}(\varepsilon)v^{j}.$ (2.34)
Finally, the LRL vector (2.4) yields
$P^{i}[A^{j}]=\frac{\partial{A^{j}}}{\partial{\dot{r}^{i}}}=2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k},\qquad{\bf X}_{A^{j}}=(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})\partial_{r^{i}}.$ (2.35)
In contrast to the previous generators, this expression is linear in $\dot{r}^{i}=v^{i}$ but not strictly linear, namely $\dfrac{\partial^{2}A^{j}}{\partial\dot{r}^{k}\partial\dot{r}^{i}}=2\delta_{ik}r_{j}-\delta_{jk}r_{i}-\delta_{ij}r_{k}$ is not proportional to $\delta_{ik}$. As a consequence, it represents the generator of a genuine dynamical symmetry. The resulting dynamical symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ involve the prolongation of this generator (2.35) to phase space. Its components consist of $P^{i}[A^{j}]$ and $\mathcal{D}_{t}P^{i}[A^{j}]$, where
$\mathcal{D}_{t}r^{k}=v^{k},\quad\mathcal{D}_{t}v^{k}=-\kappa|r|^{-1}\hat{r}^{k},\quad|r|=\sqrt{r^{k}r_{k}}.$ (2.36)
This yields
$\hat{\bf X}_{A^{j}}=\big(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k}\big)\partial_{r^{i}}+\big(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})\big)\partial_{v^{i}}.$ (2.37)
Thus, on solutions $(r^{i}(t),v^{i}(t))$, the dynamical symmetry transformations arising from the LRL vector are given by
$r^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{A^{j}})r^{i}(t),\quad v^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{A^{j}})v^{i}(t).$ (2.38)
The explicit form of these transformations will be obtained in section 5.
As pointed out in section 1, the underlying constants of motion $E$, $L^{i}$, $A^{i}$ are not all independent, since the magnitude of $A^{i}$ is a function of $E$ and $L^{i}$. The physically relevant content of $A^{i}$ is its direction:
$\Theta^{i}:=|A|^{-1}A^{i},$ (2.39)
where
$|A|=\sqrt{A^{j}A_{j}}=\sqrt{\kappa^{2}+2EL^{j}L_{j}}.$ (2.40)
This direction vector (2.39) is a constant of motion, and hence it gives rise to an infinitesimal symmetry just like $A^{i}$ does. Applying Proposition 1 yields
$P^{i}[\Theta^{j}]=\frac{\partial{\Theta^{j}}}{\partial{\dot{r}^{i}}}=|A|^{-1}(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})+2|A|^{-3}E(r^{i}\dot{r}^{k}-r^{k}\dot{r}^{i})r_{k}A^{j}$ (2.41)
using $|A|^{2}=\kappa^{2}+2EL^{k}L_{k}$ and $L^{k}L_{k}=r_{k}\dot{r}_{l}(r^{k}\dot{r}^{l}-r^{l}\dot{r}^{k})$. Observe that $r^{i}\dot{r}^{k}-r^{k}\dot{r}^{i}=\epsilon^{ikl}L_{l}$, and thus the infinitesimal symmetry generator is given by
${\bf X}_{\Theta^{j}}=\big(|A|^{-1}(2\dot{r}^{i}r^{j}-r^{i}\dot{r}^{j}-\delta^{ij}\dot{r}^{k}r_{k})+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}r_{k}\big)\partial_{r^{i}}.$ (2.42)
This represents a genuine dynamical symmetry, similarly to the previous generator (2.35).
To prolong the generator (2.42) to phase space, its component $\mathcal{D}_{t}P^{i}[\Theta^{j}]$ must be computed. Since $L_{k}$, $E$, $A_{k}$ are constants of motion, they satisfy $\mathcal{D}_{t}L_{k}=\mathcal{D}_{t}E=\mathcal{D}_{t}A_{k}=0$. Hence,
$\mathcal{D}_{t}P^{i}[\Theta^{j}]=|A|^{-1}(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}))+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}v_{k}$ (2.43)
by use of the phase-space equations of motion (2.36). Thus,
$\hat{\bf X}_{\Theta^{j}}$ $=\big(|A|^{-1}(2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k})+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}r_{k}\big)\partial_{r^{i}}$ (2.44)
$\qquad+\big(|A|^{-1}(v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}))+2|A|^{-3}E\epsilon^{ikl}L_{l}A^{j}v_{k}\big)\partial_{v^{i}}$
is the prolonged generator in phase space.
The resulting dynamical symmetry transformations on solutions $(r^{i}(t),v^{i}(t))$ are then given by
$r^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{\Theta^{j}})r^{i}(t),\quad v^{i}(t){}^{*}=\exp(\varepsilon^{j}\hat{\bf X}_{\Theta^{j}})v^{i}(t).$ (2.45)
Their explicit form will be obtained in section 4.
To conclude the present section, the infinitesimal symmetry generators ${\bf X}_{E}$, ${\bf X}_{L^{j}}$, ${\bf X}_{A^{j}}$, and ${\bf X}_{\Theta^{j}}$ will be written in vector notation, as shown in Table 1.
$\begin{gathered}\text{constant}\\ \text{of motion}\end{gathered}$ symmetry generator $\bf X$}
$E$ ${\vec{v}}\cdot\partial_{{\vec{r}}}$
$\vec{L}$ ${\vec{r}}\times\partial_{{\vec{r}}}$
$\vec{A}$ $(2{\vec{r}}\ {\vec{v}}-{\vec{v}}\ {\vec{r}}-{\vec{r}}\cdot{\vec{v}}\ \mathbf{I})\cdot\partial_{{\vec{r}}}$
$\hat{\Theta}$ $|A|^{-1}\big(2{\vec{r}}{\vec{v}}-{\vec{v}}{\vec{r}}+2E|A|^{-2}\vec{A}({\vec{r}}\times\vec{L})-{\vec{r}}\cdot{\vec{v}}\ \mathbf{I}\big)\cdot\partial_{{\vec{r}}}$
Table 1. Generators of infinitesimal symmetries arising from energy, angular momentum, LRL vector, and the LRL direction vector.
3. Poisson brackets
The Kepler problem (2.1) has the Hamiltonian formulation
$\dot{{\vec{r}}}=\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{v}}}},\quad\dot{{\vec{v}}}=-\frac{\partial{H({\vec{r}},{\vec{v}})}}{\partial{{\vec{r}}}}$ (3.1)
where
$H({\vec{r}},{\vec{v}})=\tfrac{1}{2}|{\vec{v}}|^{2}-\kappa|{\vec{r}}|^{-1}$ (3.2)
is the energy (2.2). The associated Poisson bracket is given by
$\{F,G\}=\frac{\partial{F}}{\partial{{\vec{r}}}}\cdot\frac{\partial{G}}{\partial{{\vec{v}}}}-\frac{\partial{G}}{\partial{{\vec{r}}}}\cdot\frac{\partial{F}}{\partial{{\vec{v}}}}$ (3.3)
for any functions $F({\vec{r}},{\vec{v}})$ and $G({\vec{r}},{\vec{v}})$ on phase space.
As is well known, the constants of motion — energy, angular momentum , LRL vector — form a Lie algebra under the Poisson bracket. To compute the brackets, it is simplest to work in index notation (cf Section 2), with
$E$ $=\tfrac{1}{2}v^{j}v_{j}-\kappa|r|^{-1},$ (3.4)
$L^{i}$ $=\epsilon^{ijk}r_{j}v_{k},$ (3.5)
$A^{i}$ $=\epsilon^{ijk}v_{j}L_{k}-\kappa|r|^{-1}r^{i}=(v^{j}v_{j}-\kappa|r|^{-1})r^{i}-(r^{j}v_{j})v^{i}$ (3.6)
respectively being the energy (Hamiltonian), angular momentum, and LRL vector.
Then, via using expressions (2.26), (2.27), (2.35) obtained from the infinitesimal symmetries, the Poisson brackets are readily computed to be:
$\{E,L^{i}\}$ $=\frac{\partial{E}}{\partial{r^{k}}}P^{k}[L^{i}]-\frac{\partial{L^{i}}}{\partial{r^{k}}}P^{k}[E]=\kappa|r|^{-3}r_{k}\epsilon^{ilk}r_{l}-\epsilon^{ikj}v_{j}v_{k}=0,$ (3.7)
$\{L^{i},L^{j}\}$ $=\frac{\partial{L^{i}}}{\partial{r^{k}}}P^{k}[L^{j}]-\frac{\partial{L^{j}}}{\partial{r^{k}}}P^{k}[L^{i}]=\epsilon^{li}{}_{k}v_{l}\epsilon^{kjm}r_{m}-\epsilon^{lj}{}_{k}v_{l}\epsilon^{kim}r_{m}=\epsilon^{ijk}L_{k},$ (3.8)
and
$\begin{aligned} \{E,A^{i}\}&=\frac{\partial{E}}{\partial{r^{k}}}P^{k}[A^{i}]-\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[E]\\ &=\kappa|r|^{-3}r_{k}(2v^{k}r^{i}-r^{k}v^{i}-\delta^{ki}v^{l}r_{l})-(\delta^{ik}(v^{j}v_{j}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})v_{k}\\ &=0,\end{aligned}$ (3.9)
$\begin{aligned} \{A^{i},L^{j}\}&=\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[L^{j}]-\frac{\partial{L^{j}}}{\partial{r^{k}}}P^{k}[A^{i}]\\ &=(\delta^{ik}(v^{l}v_{l}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})\epsilon^{kjm}r_{m}-\epsilon^{lj}{}_{k}v_{l}(2v^{k}r^{i}-r^{k}v^{i}-\delta^{ki}v^{m}r_{m})\\ &=\epsilon^{ijk}A_{k},\end{aligned}$ (3.10)
$\begin{aligned} \{A^{i},A^{j}\}&=\frac{\partial{A^{i}}}{\partial{r^{k}}}P^{k}[A^{j}]-\frac{\partial{A^{j}}}{\partial{r^{k}}}P^{k}[A^{i}]\\ &=(\delta^{ik}(v^{l}v_{l}-\kappa|r|^{-1})-v^{i}v^{k}+\kappa|r|^{-3}r^{i}r^{k})(2v_{k}r^{j}-r_{k}v^{j}-\delta_{k}{}^{j}v_{m}r^{m})\\ &\qquad-(\delta^{jk}(v^{l}v_{l}-\kappa|r|^{-1})-v^{j}v^{k}+\kappa|r|^{-3}r^{j}r^{k})(2v_{k}r^{i}-r_{k}v^{i}-\delta_{k}{}^{i}v_{m}r^{m})\\ &=-2E\epsilon^{ijk}L_{k}.\end{aligned}$ (3.11)
Thus, $E$, $L^{i}$, and $A^{i}$ generate a closed algebra of Poisson brackets. Its structure is well known to take the simplest form if $A^{i}$ is replaced by
$M^{i}=\sqrt{2|E|}^{-1}A^{i}$ (3.12)
assuming $E\neq 0$. Notice that $M^{i}$ is a constant of motion and has the same physical dimensions as $L^{i}$. This yields the 7-dimensional algebra
$\{E,L^{i}\}=0,\quad\{E,M^{i}\}=0,\quad\{L^{i},M^{j}\}=\epsilon^{ij}{}_{k}M^{k},$ (3.13a)
$\{L^{i},L^{j}\}=\epsilon^{ij}{}_{k}L^{k},\quad\{M^{i},M^{j}\}=-{\rm sgn}(E)\,\epsilon^{ij}{}_{k}L^{k},$ (3.13b)
which has the structure
$\mathbb{R}\oplus\mathfrak{so}(4)\quad\quad\text{ when $E<0$},$ (3.14a)
$\mathbb{R}\oplus\mathfrak{so}(3,1)\quad\text{ when $E>0$}.$ (3.14b)
A key preliminary result now is that this algebra is isomorphic to the Lie algebra formed by the infinitesimal symmetries arising from Noether’s theorem (in reverse) applied to the constants of the motion. This holds in general (see Ref.[4]) for Lagrangian dynamical systems.
Proposition 3.
The infinitesimal symmetries associated to $E$, $L^{i}$, and $M^{i}$ in the Kepler problem have a commutator structure that is isomorphic to the Poisson bracket structure of these constants of motion. Therefore, these infinitesimal symmetries form a Lie algebra (3.14).
The structure of the Lie algebra (3.14) turns out to be closely tied to the use of $M^{i}$ in place of $A^{i}$. Perhaps unsurprisingly, since $M^{i}$ and $L^{i}$ have the same physical dimensions, the Lie subalgebra that they generate cannot be factored into simpler subalgebras. So it is natural to expect that if $M^{i}$ is replaced by the LRL direction vector (2.39), which is dimensionless, then the Lie subalgebra generated by this vector $\Theta^{i}$ and $L^{i}$ should have a simpler structure.
The brackets involving $\Theta^{i}$ can be computed directly from the previous brackets (3.7)–(3.11) by using the relation $A^{i}=|A|\,\Theta^{i}$ together with the property that the Poisson bracket obeys the Leibniz rule. In particular, for any function $F(r^{k},v^{k})$ on phase space, note that its bracket with $\Theta^{i}$ can be obtained from $\{F,A^{i}\}=|A|\ \{F,\Theta^{i}\}+\Theta^{i}\ \{F,|A|\}$, where $\{F,|A|\}=|A|^{-1}A_{j}\ \{F,A^{j}\}$. Hence, firstly,
$\{E,|A|\}$ $=|A|^{-1}A_{j}\ \{E,A^{j}\}=0,$ (3.15)
$\{L^{i},|A|\}$ $=|A|^{-1}A_{j}\ \{L^{i},A^{j}\}=|A|^{-1}\epsilon^{ijk}A_{j}A_{k}=0,$ (3.16)
$\{A^{i},|A|\}$ $=|A|^{-1}A_{j}\ \{A^{i},A^{j}\}=-2|A|^{-1}E\epsilon^{ijk}A_{j}L_{k};$ (3.17)
and then secondly, $\{F,\Theta^{i}\}=|A|^{-1}(\{F,A^{i}\}-\Theta^{i}\ \{F,|A|\})$ yields
$\{E,\Theta^{i}\}=0,$ (3.18)
$\{L^{j},\Theta^{i}\}=|A|^{-1}\epsilon^{jik}A_{k}=\epsilon^{jik}\Theta_{k},$ (3.19)
$\{|A|,\Theta^{i}\}=2|A|^{-1}E\epsilon^{ijk}A_{j}L_{k},$ (3.20)
$\{A^{j},\Theta^{i}\}=2|A|^{-1}E(\epsilon^{jlk}\Theta_{l}\Theta^{i}-\epsilon^{jik})L_{k},$ (3.21)
$\begin{aligned} \{\Theta^{j},\Theta^{i}\}&=-2|A|^{-2}E(\epsilon^{jik}-(\epsilon^{jlk}\Theta^{i}-\epsilon^{ilk}\Theta^{j})\Theta_{l}))L_{k}\\ &=-2|A|^{-2}E\epsilon^{jik}((1-\Theta_{l}\Theta^{l})L_{k}+\Theta^{l}L_{l}\Theta_{k})=0,\end{aligned}$ (3.22)
where the last line has used the vector identity $\epsilon^{jlk}F^{i}-\epsilon^{ilk}F^{j}=\epsilon^{jik}F^{l}-\epsilon^{jil}F^{k}$, as well as the relations $\Theta^{j}L_{j}=0$ and $\Theta^{j}\Theta_{j}=1$.
Therefore, only the bracket (3.19) among $E$, $L^{i}$, $\Theta^{i}$ is non-vanishing. This bracket shows that $L^{j}$ acts as an infinitesimal rotation on the direction vector $\Theta^{i}$:
$\{L^{i},\Theta^{j}\}=\epsilon^{ij}{}_{k}\Theta^{k}.$ (3.23)
Theorem 1.
The constants of motion $E$, $L^{i}$, and $\Theta^{i}$ generate a 7-dimensional Lie algebra under the Poisson bracket, which has the structure
$\mathbb{R}\oplus\mathfrak{so}(3)\rtimes\mathbb{R}^{3}.$ (3.24)
This algebra is isomorphic to the Lie algebra of the infinitesimal symmetries associated to these constants of motion.
3.1. Symmetry action on constants of motion
The set of all constants of motion of the Kepler problem forms a vector space. Any infinitesimal symmetry $\hat{\bf X}$, acting on phase space, maps this vector space into itself. The mapping is given by the Poisson bracket, as shown by a direct computation.
Proposition 4.
For any constant of motion $C(r^{i},v^{i})$ in the Kepler problem, the associated infinitesimal symmetry generator ${\rm pr}{\bf X}_{C}$ acts on a function $f(r^{i},v^{i})$ on phase space by
${\rm pr}{\bf X}_{C}f=\{f,C\}$ (3.25)
modulo the equations of motion (2.1).
This yields the following result: the only non-zero actions of the infinitesimal symmetries $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{A^{j}}$, $\hat{\bf X}_{\Theta^{j}}$ on the constants of motion $E$, $L^{i}$, $A^{i}$, and $\Theta^{i}$ are given by
${\rm pr}\hat{\bf X}_{L^{j}}(L^{i})=\epsilon^{ij}{}_{k}L^{k},$ (3.26)
${\rm pr}\hat{\bf X}_{A^{j}}(A^{i})=-2E\epsilon^{ij}{}_{k}L^{k},$ (3.27)
${\rm pr}\hat{\bf X}_{L^{j}}(A^{i})=-{\rm pr}\hat{\bf X}_{A^{i}}(L^{j})=\epsilon^{ij}{}_{k}A^{k},$ (3.28)
${\rm pr}\hat{\bf X}_{L^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(L^{j})=\epsilon^{ij}{}_{k}\Theta^{k},$ (3.29)
${\rm pr}\hat{\bf X}_{A^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(A^{j})=2|A|^{-1}E(\epsilon^{jik}-\epsilon^{jlk}\Theta_{l}\Theta^{i})L_{k}.$ (3.30)
Additionally, Proposition 4 applied to the brackets (3.13) shows that the 7-dimensional vector space spanned by $E$, $L^{i}$, $M^{i}$ carries a homomorphic representation of the LRL symmetry algebra (3.14). The same statement holds for the symmetry algebra (3.24) and the 7-dimensional vector space spanned by $E$, $L^{i}$, $\Theta^{i}$.
3.2. Dynamical symmetry algebra
Some further aspects of the correspondence between the algebra of constants of motion and the Lie algebra of infinitesimal symmetries will now be discussed.
First, quadratic invariants and Casimirs will be considered. For a given Lie algebra of Poisson brackets on phase space, a quadratic scalar $I$ is invariant if it Poisson commutes with all of the constants of motion $C$ in the Lie algebra. Replacement of the constants of motion in $I$ by their corresponding symmetry generators yields a Casimir element in the enveloping algebra of all symmetry generators.
For the Lie algebra (3.14) generated by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{M^{j}}$, it is well known that [13] the quadratic invariants consist of (linear combinations of)
$E^{2},\quad M^{j}M_{j}-{\rm sgn}(E)\,L^{j}L_{j}$ (3.31)
which can be verified by a simple computation via the Poisson brackets (3.13).
For the Lie algebra generated by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{\Theta^{j}}$, there is just the single quadratic invariant $E^{2}$, as seen from the only non-vanishing Poisson bracket (3.24).
Any invariant is manifestly a constant of motion in the Kepler problem. More generally, any function of the 7 basic constants of motion $E$, $L^{i}$, $A^{i}$ is again a constant of motion. This means that the set of all constants of motion is infinite dimensional. As a consequence, the set of all infinitesimal symmetry generators, arising from Noether’s theorem in reverse, forms an infinite-dimensional Lie algebra. In this Lie algebra, the infinitesimal generators of point symmetries comprise a finite dimensional Lie subalgebra, which is generated by $\hat{\bf X}_{E}$ and $\hat{\bf X}_{L^{j}}$, namely time-translations and rotations. Quotienting out this subalgebra leaves an infinite-dimensional Lie algebra comprised of genuine infinitesimal dynamical symmetries.
Thus, the finite-dimensional Lie algebras generated respectively by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{M^{j}}$, and by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{\Theta^{j}}$, are particular subalgebras in the infinite-dimensional Lie algebra of genuine infinitesimal dynamical symmetries in the Kepler problem.
4. Symmetry group generated by the directional LRL vector
The dynamical symmetry transformations (2.45) arising from the LRL direction vector (2.39) are most readily expressed in an explicit form by working in the coordinate space $(t,{\vec{r}},{\vec{v}})$ where the generator of the infinitesimal symmetry possesses gauge freedom which can be exploited to simplify the transformations. This generator has the form
${\bf Y}_{\Theta^{j}}=\tau^{j}D_{t}+\hat{\bf X}_{\Theta^{j}}$ (4.1)
where $\hat{\bf X}_{\Theta^{j}}$ is the generator (2.44) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. In component form,
${\bf Y}_{\Theta^{j}}(t)$ $=\tau^{j},$ (4.2a)
${\bf Y}_{\Theta^{j}}(r^{i})$ $=\tau^{j}v^{i}+2v^{i}r^{j}-v^{j}r^{i}-\delta^{ij}v^{k}r_{k}+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}r_{k},$ (4.2b)
${\bf Y}_{\Theta^{j}}(v^{i})$ $=-\kappa|r|^{-3}\tau^{j}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}v_{k}.$ (4.2c)
Now $\tau^{j}$ can be chosen so that one component of this generator is zero, which will consequently be invariant under the infinitesimal symmetry (4.1).
It will be useful to have $|r|$ be invariant, ${\bf Y}_{\Theta^{j}}(|r|)=0$. This condition is easily computed via
${\bf Y}_{\Theta^{j}}(|r|^{2})$ $=2|r|{\bf Y}_{\Theta^{j}}(|r|)$ (4.3)
$=2r_{i}{\bf Y}_{\Theta^{j}}(r^{i})=(\tau^{j}+r^{j})v^{i}r_{i}-v^{j}r^{i}r_{i},$
which yields
$\tau^{j}=(v^{j}r^{i}-r^{j}v^{i})r_{i}=-\epsilon^{jik}r_{i}L_{k}.$ (4.4)
Hence, this fixes all of the gauge freedom.
The components of the gauge-fixed generator are given by
${\bf Y}_{\Theta^{j}}(t)$ $=-\epsilon^{jik}L_{k}r_{i},$ (4.5a)
${\bf Y}_{\Theta^{j}}(r^{i})$ $=-\epsilon^{jlk}r_{l}L_{k}v^{i}+2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k}+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}r_{k},$ (4.5b)
${\bf Y}_{\Theta^{j}}(v^{i})$ $=\kappa|r|^{-3}\epsilon^{jlk}r_{l}L_{k}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}v_{k}.$ (4.5c)
This generator represents a vector field in the coordinate space $(t,r^{i},v^{i})$, and its flow thereby defines the dynamical symmetry transformations
$(t,r^{i},v^{i})\to(t,r^{i},v^{i})^{*}=\exp(\varepsilon^{j}{\bf Y}_{\Theta^{j}})(t,r^{i},v^{i})$ (4.6)
acting in this space, where $\varepsilon^{j}$ is the parameter of the flow.
The flow (4.6) satisfies the system of 7 coupled differential equations
$\frac{d{t^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}[\varepsilon](t^{*}(\varepsilon)),\quad\frac{d{r^{i}{}^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}[\varepsilon](r^{i}{}^{*}(\varepsilon)),\quad\frac{d{v^{i}{}^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}[\varepsilon](v^{i}{}^{*}(\varepsilon))$ (4.7)
where ${\bf Y}_{\Theta^{j}}^{*}[\varepsilon]={\bf Y}_{\Theta^{j}}\big|_{t=(t^{*}(\varepsilon),r^{i}=r^{i}{}^{*}(\varepsilon),v^{i}=v^{i}{}^{*}(\varepsilon)}$ denotes the generator evaluated on the flow. In addition, the flow satisfies the condition that $\varepsilon^{j}=0$ gives the identity transformation,
$t^{*}(0)=t,\quad r^{i}{}^{*}(0)=r^{i},\quad v^{i}{}^{*}(0)=v^{i}.$ (4.8)
The goal is to integrate this system explicitly. Hereafter, for notational simplicity, the dependence on $\varepsilon^{j}$ will be suppressed in the transformed variables and the generator.
As a starting point, the chosen gauge condition implies that
$\frac{d{|r|^{*}}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}(|r|^{*})=0,$ (4.9)
and therefore,
$|r|^{*}=|r|.$ (4.10)
This constitutes the solution of one combination of the 7 differential equations in the system (4.7).
The strategy for integrating the remainder of the system will involve the following three steps.
(1) The first step will be to convert the differential equations for $r^{i}{}^{*}$ and $v^{i}{}^{*}$ into an equivalent simpler system with the constants of motion $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$ taken as the variables in the flow, together with the invariant $|r|$. This system is obtained from the action of the infinitesimal symmetry $\hat{\bf X}_{\Theta^{i}}$ on $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, which is given by the Poisson brackets.
(2) The second step will consist of expressing $r^{i}{}^{*}$ and $v^{i}{}^{*}$ algebraically in terms of $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, and $|r|$. This amounts to a straightforward linear algebra computation.
(3) The third step will be to directly integrate the remaining differential equation for $t{}^{*}$, as its righthand side will involve only solved quantities.
Details are carried out in the appendix.
Theorem 2.
The dynamical symmetry transformations (4.6), which arise from the LRL direction vector (2.39), have the explicit form
$t{}^{*}=t+\int_{0}^{\vec{\varepsilon}}\big(\alpha^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}+\beta^{r}{}^{*}|\vec{L}^{*}|^{2}\vec{\Theta}\big)\cdot d\vec{\varepsilon},$ (4.11a)
${\vec{r}}{}^{*}=\alpha^{r}{}^{*}\vec{\Theta}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta},\quad{\vec{v}}{}^{*}=\alpha^{v}{}^{*}\vec{\Theta}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta}$ (4.11b)
with parameter $\vec{\varepsilon}$, where
$\alpha^{r}{}^{*}=|\vec{L}^{*}|^{2}-\kappa|{\vec{r}}|,\quad\beta^{r}{}^{*}={\rm sgn}({\vec{r}}\cdot{\vec{v}})|{\vec{r}}|\sqrt{2(E+\kappa|{\vec{r}}|^{-1})-(|{\vec{r}}|^{-1}|\vec{L}{\mathstrut}^{*}|)^{2}},$ (4.12a)
$\alpha^{v}{}^{*}=-{\rm sgn}({\vec{r}}\cdot{\vec{v}})\kappa\sqrt{2(E+\kappa|{\vec{r}}|^{-1})-(|{\vec{r}}|^{-1}|\vec{L}{\mathstrut}^{*}|)^{2}},\quad\beta^{v}{}^{*}=2E+\kappa|{\vec{r}}|^{-1},$ (4.12b)
and
$\vec{L}^{*}=\vec{L}+\vec{\varepsilon}\times\vec{\Theta}.$ (4.13)
Under these transformations, $E$ and $\vec{\Theta}$ are invariant, along with $|{\vec{r}}|$.
As shown by Theorem 1, the three infinitesimal symmetries $\hat{\bf X}_{\Theta^{j}}$ form an abelian Lie algebra $\mathbb{R}^{3}$. This implies that the set of all dynamical symmetry transformations (4.11) under composition generates a 3-dimensional abelian Lie group.
The physical meaning of this symmetry group is that it preserves energy and the direction of the LRL vector, while angular momentum is shifted by a vector in the plane orthogonal to the LRL vector. Hence, the magnitude of the LRL vector is changed:
$|\vec{A}^{*}|=\sqrt{|\vec{A}|^{2}+2E\big(\vec{\varepsilon}\cdot(\vec{\Theta}\times\vec{L})+|\vec{\varepsilon}|^{2}-(\vec{\varepsilon}\cdot\vec{\Theta})^{2}\big)}.$ (4.14)
This change is proportional to the change in eccentricity. In the case of elliptical orbits, since energy is unchanged, the period $T=2\pi/\sqrt{-2E}$ and semi-major axis $a=\kappa/(-2E)$ are also preserved.
5. LRL symmetry transformation group
The explicit form of the LRL dynamical symmetry transformations (2.38) will now be derived in a similar way to the result in Theorem 2.
In the coordinate space $(t,{\vec{r}},{\vec{v}})$, the LRL infinitesimal symmetry generator is given by
${\bf Y}_{A^{j}}=\tau^{j}D_{t}+\hat{\bf X}_{A^{j}}$ (5.1)
where $\hat{\bf X}_{A^{j}}$ is the symmetry generator (2.37) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. This gauge freedom will be fixed by taking $|r|$ to be invariant:
${\bf Y}_{A^{j}}(|r|^{2})$ $=2|r|{\bf Y}_{A^{j}}(|r|)$ (5.2)
$=2r_{i}{\bf Y}_{A^{j}}(r^{i})=(\tau^{j}+r^{j})v^{i}r_{i}-v^{j}r^{i}r_{i},$
which yields expression (4.4) for $\tau^{j}$. The gauge-fixed generator thus has the components
${\bf Y}_{A^{j}}(t)$ $=-\epsilon^{jik}L_{k}r_{i},$ (5.3a)
${\bf Y}_{A^{j}}(r^{i})$ $=-\epsilon^{jlk}r_{l}L_{k}v^{i}+2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k},$ (5.3b)
${\bf Y}_{A^{j}}(v^{i})$ $=\kappa|r|^{-3}\epsilon^{jlk}r_{l}L_{k}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}),$ (5.3c)
which define a vector field in the coordinate space $(t,r^{i},v^{i})$.
The flow of this vector field determines the LRL dynamical symmetry transformations
$(t,r^{i},v^{i})\to(t,r^{i},v^{i})^{*}=\exp(\varepsilon^{j}{\bf Y}_{A^{j}})(t,r^{i},v^{i})$ (5.4)
where $\varepsilon^{j}$ is the parameter of the flow.
In component form, the flow is defined by a system of 7 coupled differential equations
$\frac{d{t^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{A^{j}}^{*}[\varepsilon](t^{*}(\varepsilon)),\quad\frac{d{r^{i}{}^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{A^{j}}^{*}[\varepsilon](r^{i}{}^{*}(\varepsilon)),\quad\frac{d{v^{i}{}^{*}(\varepsilon)}}{d{\varepsilon_{j}}}={\bf Y}_{A^{j}}^{*}[\varepsilon](v^{i}{}^{*}(\varepsilon))$ (5.5)
where ${\bf Y}_{A^{j}}^{*}[\varepsilon]={\bf Y}_{A^{j}}\big|_{t=(t^{*}(\varepsilon),r^{i}=r^{i}{}^{*}(\varepsilon),v^{i}=v^{i}{}^{*}(\varepsilon)}$ denotes the generator evaluated on the flow. The flow also satisfies the condition (4.8) stating that $\varepsilon^{j}=0$ gives the identity transformation. For notational simplicity, hereafter the dependence on $\varepsilon^{j}$ will be suppressed.
To integrate this system, the same three steps used in the previous section will be employed with one change: In the first step, the primary variables will be taken to be $E^{*}$ and $|r|$, together with $M^{i}{}^{*}$ and $L^{i}{}^{*}$, as the latter satisfy a coupled system shown by their Poisson brackets (3.13), with $\Theta^{i}{}^{*}$ no longer being invariant in the flow. This leads to a more complicated transformation on $r^{i}{}^{*}$ and $v^{i}{}^{*}$ in the second step. Details of all three steps are carried out in the appendix.
To state the main result, introduce the matrix-rotation operator
$\underline{R}(\phi,\hat{n})=\cos(\phi)\ \underline{I}+(1-\cos(\phi))\underline{I}\,\hat{n}\,\hat{n}\cdot+\sin(\phi)\underline{S}\,\hat{\varepsilon}\times$ (5.6)
with
$\underline{I}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad\underline{S}=\sqrt{-{\rm sgn}(E)}\begin{pmatrix}0&1\\ {\rm sgn}(E)&0\end{pmatrix}$ (5.7)
where $\phi$ is a scalar parameter and $\hat{n}$ is a unit-vector parameter. For $E<0$, this operator is a twisted composition of an internal rotation $SO(2)$ and a spatial rotation $SO(3)$. For $E>0$, the internal part becomes a hyperbolic rotation $SO(1,1)$.
Theorem 3.
The LRL dynamical symmetry transformations (5.4) have the explicit form
$t{}^{*}=t+\int_{0}^{\vec{\varepsilon}}\big(\alpha^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*}+\beta^{r}{}^{*}|\vec{L}^{*}|^{2}\vec{\Theta}^{*}\big)\cdot d\vec{\varepsilon},$ (5.8a)
${\vec{r}}{}^{*}=\alpha^{r}{}^{*}\vec{\Theta}^{*}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*},\quad{\vec{v}}{}^{*}=\alpha^{v}{}^{*}\vec{\Theta}^{*}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*}$ (5.8b)
with parameter $\vec{\varepsilon}$, where the coefficients are given by expressions (4.12).
In the case $E\neq 0$, the transformed constants of motion $\vec{L}^{*}$ and $\vec{\Theta}^{*}$ are given by
$\begin{pmatrix}\vec{L}^{*}\\ \vec{M}^{*}\end{pmatrix}=\underline{R}(\sqrt{-2E}\,|\vec{\varepsilon}|,\hat{\varepsilon})\begin{pmatrix}\vec{L}\\ \vec{M}\end{pmatrix},\quad\hat{\varepsilon}=|\vec{\varepsilon}|^{-1}\vec{\varepsilon},$ (5.9)
in terms of $\vec{M}=|\vec{M}|\vec{\Theta}$ with $|\vec{M}|^{2}={\rm sgn}(E)|\vec{L}|^{2}+\tfrac{1}{2}\kappa^{2}|E|^{-1}$. Under these transformations, $E$ and $|\vec{M}|^{2}-{\rm sgn}(E)\,|\vec{L}|^{2}$, along with $|{\vec{r}}|$, are invariant.
In the case $E=0$,
$\vec{\Theta}{}^{*}=\vec{\Theta},\quad\vec{L}^{*}=\vec{L}+|\vec{A}|\,\vec{\varepsilon}\times\vec{\Theta}$ (5.10)
are the transformed constants of motion, with $|A|^{2}=\kappa^{2}+2E|\vec{L}|^{2}$. Under these transformations, $E$, $\vec{\Theta}$, and $|{\vec{r}}|$ are invariant, while the magnitude of the LRL vector is changed:
$|\vec{A}^{*}|=|\vec{A}|\sqrt{1+2E|\vec{A}|{\mathstrut}^{-1}\big(\vec{\varepsilon}\cdot(\vec{\Theta}\times\vec{L})+|\vec{\varepsilon}|^{2}-(\vec{\varepsilon}\cdot\vec{\Theta})^{2}\big)}.$ (5.11)
From Proposition 3, note that the three infinitesimal symmetries $\hat{\bf X}_{A^{j}}$ do not comprise a closed algebra, but instead they generate a 6-dimensional Lie algebra that contains the spatial rotations $\mathfrak{so}(3)$ as a subalgebra generated by $\hat{\bf X}_{L^{j}}$ through the commutator of $\hat{\bf X}_{A^{j}},\hat{\bf X}_{A^{k}}]$. Consequently, the LRL dynamical symmetry transformations (5.8) do not themselves form a group. However, if the parameter $\hat{\varepsilon}$ is taken to be fixed, then the set of resulting transformations generates a one-dimensional Lie group with parameter $\phi=\sqrt{-2E}\,|\vec{\varepsilon}|$. For $E<0$, the group is isomorphic to the rotation group $SO(2)$, while for $E>0$, it is isomorphic to the hyperbolic rotation group $SO(1,1)$. When $E=0$, the group is isomorphic to $\mathbb{R}$. In each case, the composition of all these one-dimensional groups will generate a 6-dimensional Lie group whose structure depends on $E$, which is $SO(4)$ when $E<0$; $SO(3,1)$ when $E<0$; $SO(3)\rtimes\mathbb{R}^{3}$ when $E=0$.
This symmetry transformation group acts as isometries in the space of constants of motion $(\vec{L},\vec{M})$, where $\vec{L}\cdot\vec{M}=0$ and $|\vec{L}|^{2}-{\rm sgn}(E)|\vec{M}|^{2}=\kappa^{2}/(2|E|)$, with $|\vec{L}|\neq 0$ and $|\vec{M}|\neq 0$. When $E<0$, the space is isomorphic to $S^{2}\times S^{1}\times S^{1}$, which has the symmetry group $SO(4)$. The space is isomorphic to $S^{2}\times S^{1}\times\mathbb{R}$ when $E>0$, and the symmetry group is then $SO(3,1)$.
The LRL dynamical symmetry transformations (5.8) have the physical meaning that they preserve energy, while angular momentum and the LRL vector are changed in a specific way that depends on the energy. In the case $E<0$, the vectors $\vec{L}$ and $\vec{M}=\sqrt{2|E|}^{-1}\vec{A}$ are rotated, preserving $|\vec{L}|^{2}+|\vec{M}|^{2}$ and $\vec{L}\cdot\vec{M}=0$, along with the period $T=2\pi/\sqrt{2|E|}$ and semi-major axis $a=\kappa/(2|E|)$ of the elliptical orbit. Likewise in the case $E>0$, $|\vec{L}|^{2}-\frac{1}{2}|\vec{A}|^{2}/E$ is preserved, as is the semi-major axis $a=\kappa/(2E)$ of the hyperbolic orbit. In the case $E=0$, $\vec{A}$ is preserved and $\vec{L}$ is shifted by a vector in the plane orthogonal to $\vec{A}$.
6. Concluding remarks
In the Kepler problem, the dynamical symmetry transformations arising respectively from the LRL vector and the LRL direction vector through Noether’s theorem (in reverse) have been derived in an explicit form in terms of the kinematic variables $t$, $\vec{r}$}, $\vec{v}$}. The structure and physical meaning of the symmetry groups generated by these respective transformations (5.8) and (4.11) have been explained.
There are equivalent symmetry transformations that act in phase space, with the time $t$ being invariant. Their simplest formulation is given by considering the set of solutions $({\vec{r}}(t),{\vec{v}}(t))$ of the equations of motion (2.1). Under either of the dynamical symmetry transformations, each solution is mapped into a family of solutions
$({\vec{r}}(t),{\vec{v}}(t))\to({\vec{r}}^{\dagger}(t;\vec{\varepsilon}),{\vec{v}}^{\dagger}(t;\vec{\varepsilon}))$ (6.1)
with parameter $\vec{\varepsilon}$. To write the mapping in terms of the symmetry transformations, it is helpful to indicate the explicit dependence of these transformations (5.8) and (4.11) on the kinematic variables and the parameter: ${\vec{r}}^{*}={\vec{r}}^{*}[r,v;\varepsilon]$, ${\vec{v}}^{*}={\vec{v}}^{*}[r,v;\varepsilon]$, $t^{*}=t^{*}[t,r,v;\varepsilon]$. Now, for a fixed parameter $\vec{\epsilon}$, each solution represents a curve whose points $(t,{\vec{r}}(t),{\vec{v}}(t))$ get mapped to points with coordinates $\big(t^{*}[t,r(t),v(t);\varepsilon],{\vec{r}}^{*}[r(t),v(t);\varepsilon],{\vec{v}}^{*}[r(t),v(t);\varepsilon]\big)$ which lie on the curve $({\vec{r}}^{\dagger}(t;\vec{\varepsilon}),{\vec{v}}^{\dagger}(t;\vec{\varepsilon}))$, as shown in Fig. 6. This implies the relation
${\vec{r}}^{*}[r^{\dagger},v^{\dagger};-\varepsilon]={\vec{r}}(t^{*}(t,r^{\dagger},v^{\dagger};-\varepsilon)),\quad{\vec{v}}^{*}[r^{\dagger},v^{\dagger};-\varepsilon]={\vec{v}}(t^{*}(t,r^{\dagger},v^{\dagger};-\varepsilon))$ (6.2)
giving an implicit system of equations through which the coordinate values $({\vec{r}}^{\dagger},{\vec{v}}^{\dagger})$ can be determined at each time $t$. Hence, an implicit formulation of the mapping (6.1) is obtained.
[…]
All of the methods and results in the present work can be extended, firstly, to general central force dynamics. In particular, a generalization of the conserved LRL vector is known to exist for motion under any central force [13, 14, 15, 16, 17, 11] in $n>1$ dimensions. The physically relevant part of this vector is its direction in the plane of motion, which will give rise to an associated infinitesimal symmetry through Noether’s theorem and hence generate a corresponding transformation group of dynamical symmetries. A sequel paper [18] will study this group as well as the group containing the generalized LRL symmetry transformations.
A further extension can be contemplated to special relativity (see e.g. Ref.[13]). More interestingly, as shown recently in Ref.[19], an analog of the LRL vector exists for particle orbits in Schwarzschild spacetime, where these orbits can be viewed as motion under a relativistic, gravitational central force [20]. Finding the transformation group of dynamical symmetries that is generated by this analog is an interesting question which will be pursued in future work [21].
Appendix A Proof of Theorem 2
(1) The constants of motion $E$, $L^{i}$, $\Theta^{i}$ constitute 5 independent quantities, since they satisfy the two relations $L^{i}\Theta_{i}=0$ and $\Theta^{i}\Theta_{i}=1$. Together with $|r|$, they represent 6 degrees of freedom.
For any constant of motion $C$ in the Kepler problem, observe that
${\bf Y}_{\Theta^{j}}C=\tau^{j}D_{t}C+\hat{\bf X}_{\Theta^{j}}(C)=\hat{\bf X}_{\Theta^{j}}(C)$ (A.1)
holds on solutions of the equation of motion (2.1), since $D_{t}C=\mathcal{D}_{t}C=0$. This yields a differential equation
$\frac{d{C^{*}}}{d{\varepsilon_{j}}}=\hat{\bf X}_{\Theta^{j}}^{*}(C^{*})$ (A.2)
for the evolution of $C^{*}$ under the flow.
Use of expression (3.29) for the non-vanishing action of $\hat{\bf X}_{\Theta^{j}}$ on $E$, $L^{i}$, $\Theta^{i}$ yields a system of simple differential equations
$\frac{d{E^{*}}}{d{\varepsilon_{j}}}=0,\quad\frac{d{\Theta^{i}{}^{*}}}{d{\varepsilon_{j}}}=0,\quad\frac{d{L^{i}{}^{*}}}{d{\varepsilon_{j}}}=\epsilon^{ij}{}_{k}\Theta^{k}{}^{*}.$ (A.3)
Their solution is given by
$E^{*}=E,\quad\Theta^{i}{}^{*}=\Theta^{i},\quad L^{i}{}^{*}=\epsilon^{i}{}_{jk}\varepsilon^{j}\Theta^{k}+L^{i}.$ (A.4)
(2) Since $r^{i}$ and $v^{i}$ represent vectors that lie in the plane of motion and are linearly independent, they can be expressed as linear combinations of $\Theta^{i}$ and $\epsilon^{i}{}_{jk}L^{j}\Theta^{k}$ which are a pair of orthogonal vectors lying in the plane of motion. Their explicit form is given by
$\Theta^{i}$ $=\epsilon^{ijk}v_{j}L_{k}-\kappa|r|^{-1}r^{i}=(v^{j}v_{j}-\kappa|r|^{-1})r^{i}-(r^{j}v_{j})v^{i},$ (A.5)
$\epsilon^{i}{}_{jk}L^{j}\Theta^{k}$ $=v^{i}L^{j}L_{j}-\kappa|r|^{-1}\epsilon^{ijk}L_{j}r_{k}=(|r|^{2}(v^{j}v_{j}-\kappa|r|^{-1})-(r^{j}v_{j})^{2})v^{i}+(\kappa|r|^{-1}r^{j}v_{j})r^{i}.$ (A.6)
Thus,
$r^{i}=\alpha^{r}\Theta^{i}+\beta^{r}\epsilon^{i}{}_{kl}L^{k}\Theta^{l},\quad v^{i}=\alpha^{v}\Theta^{i}+\beta^{v}\epsilon^{i}{}_{kl}L^{k}\Theta^{l},$ (A.7)
where the coefficients in this basis expansion are given by
$\alpha^{r}=r_{j}\Theta^{j}=|L|^{2}-\kappa|r|,$ $\beta^{r}=|L|^{-2}r_{j}\epsilon^{j}{}_{kl}L^{k}\Theta^{l}=r^{j}v_{j},$ (A.8a)
$\alpha^{v}=v_{j}\Theta^{j}=-\kappa|r|^{-1}r^{j}v_{j},$ $\beta^{v}=|L|^{-2}v_{j}\epsilon^{j}{}_{kl}L^{k}\Theta^{l}=v^{j}v_{j}-\kappa|r|^{-1},$ (A.8b)
with
$|L|^{2}=L^{j}L_{j}=|r|^{2}v^{j}v_{j}-(r^{j}v_{j})^{2}.$ (A.9)
Now, via the expression (3.4) for energy, note that
$v^{j}v_{j}=2(E+\kappa|r|^{-1}),\quad r^{j}v_{j}=\pm\sqrt{|r|^{2}v^{j}v_{j}-|L|^{2}}=\pm\sqrt{2|r|(|r|E+\kappa)-|L|^{2}}$ (A.10)
where the $\pm$ denotes the sign of the radial velocity, ${\rm sgn}({\vec{r}}\cdot{\vec{v}})$. Substituting these expressions into the coefficients (A.8) then yields
$\alpha^{r}=|L|^{2}-\kappa|r|,\quad\beta^{r}={\rm sgn}({\vec{r}}\cdot{\vec{v}})|r|\sqrt{2(E+\kappa|r|^{-1})-(|r|^{-1}|L|)^{2}},$ (A.11a)
$\alpha^{v}=-{\rm sgn}({\vec{r}}\cdot{\vec{v}})\kappa\sqrt{2(E+\kappa|r|^{-1})-(|r|^{-1}|L|)^{2}},\quad\beta^{v}=2E+\kappa|r|^{-1},$ (A.11b)
which are expressed solely in terms of the constants of motion $E$ and $|L|$ and the invariant $|r|$.
Finally, under the flow (A.4), the expansion (A.7) yields the transformations (4.11b) on $r^{i}{}^{*}$ and $v^{i}{}^{*}$.
(3) The flow equation for $t^{*}$ now becomes
$\frac{d{t^{*}}}{d{\varepsilon_{j}}}=-\epsilon^{jik}r_{i}{}^{*}L_{k}{}^{*},$ (A.12)
which can be directly integrated as a quadrature. Substituting the transformation for for $r^{i}{}^{*}$ gives the transformation (4.11a).
Appendix B Proof of Theorem 3
(1) The 7 constants of motion $E$, $L^{i}$, $A^{i}$, together with $|r|$, represent 6 degrees of freedom, since $L^{i}A_{i}=0$ and $|A|^{2}=\kappa^{2}+2EL^{i}L_{i}$ hold identically. Their evolution under the flow (5.4) is given by the differential equation
$\frac{d{C^{*}}}{d{\varepsilon_{j}}}=\hat{\bf X}_{A^{j}}^{*}(C^{*}).$ (B.1)
Use of expressions (3.27)–(3.28) for the non-vanishing action of $\hat{\bf X}_{A^{j}}$ on $E$, $L^{i}$, and $A^{i}$ yields the system of 7 coupled differential equations
$\frac{d{E^{*}}}{d{\varepsilon_{j}}}=0,$ (B.2)
$\frac{d{L^{i}{}^{*}}}{d{\varepsilon_{j}}}=\epsilon^{ij}{}_{k}A^{k}{}^{*},\quad\frac{d{A^{i}{}^{*}}}{d{\varepsilon_{j}}}=-2E^{*}\epsilon^{ij}{}_{k}L^{k}{}^{*}.$ (B.3)
Observe that the solution of the first equation (B.2) is simply
$E^{*}=E.$ (B.4)
Then the remaining coupled equations (B.3) have a fairly simple form after the change of variable (3.12):
$\dfrac{d}{d\varepsilon^{j}}\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\sqrt{2|E|}\epsilon^{ij}{}_{k}\begin{pmatrix}M^{k}{}^{*}\\ -{\rm sgn}(E)\,L^{k}{}^{*}\end{pmatrix}$ (B.5)
assuming $E\neq 0$.
In the case $E<0$, this system (B.5) is recognized to be a flow consisting of the twisted composition of an internal rotation with angular parameter $\sqrt{2|E|}$ and a spatial rotation through an angle
$|\varepsilon|=\varepsilon^{j}\varepsilon_{j},$ (B.6)
around an axis
$\hat{\varepsilon}^{j}=|\varepsilon|^{-1}\varepsilon^{j}.$ (B.7)
These two different rotations commute. The solution is given by
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\exp\big(\sqrt{2|E|}\,\underline{J}\,\varepsilon^{j}\epsilon^{i}{}_{jk}\big)\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix},\quad\underline{J}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$. (B.8)
Expanding this exponential into a Taylor series and grouping the terms into even and odd powers of $\varepsilon^{j}$ yields the matrix operator
[…]
where
$\underline{R}^{i}{}_{k}(\phi,\hat{\varepsilon}^{j})=\cos(\phi)\ \underline{I}\,\delta^{i}{}_{k}+(1-\cos(\phi))\underline{I}\,\hat{\varepsilon}^{i}\hat{\varepsilon}_{k}+\sin(\phi)\underline{J}\,\hat{\varepsilon}_{j}\epsilon^{ji}{}_{k},\quad\underline{I}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ (B.10)
represents a twisted rotation operator. Its parameters satisfy $\phi\in[0,2\pi)$, $\hat{\varepsilon}\in S^{2}$, where $S^{2}$ denotes the unit 2-sphere in $\mathbb{R}^{3}$. Hence,
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\underline{R}^{i}{}_{k}(\sqrt{2|E|}\,|\varepsilon|,\hat{\varepsilon}^{j})\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix}$. (B.11)
In the case $E>0$, the flow (B.5) consists of the twisted composition of the preceding spatial rotation with angle $|\varepsilon|$ and axis $\hat{\varepsilon}^{j}$, and an internal hyperbolic rotation (boost) with parameter $\sqrt{2|E|}$. These two different rotations again commute. Thus, the solution is given by
$\begin{pmatrix}L^{i}{}^{*}\\ M^{i}{}^{*}\end{pmatrix}=\underline{\widetilde{R}}^{i}{}_{k}(\sqrt{2|E|}\,|\varepsilon|,\hat{\varepsilon}^{j})\begin{pmatrix}L^{k}\\ M^{k}\end{pmatrix}$ (B.12)
where
$\underline{\widetilde{R}}^{i}{}_{k}(\psi,\hat{\varepsilon}^{j})=\cosh(\psi)\ \underline{I}\,\delta^{i}{}_{k}+(1-\cosh(\psi))\underline{I}\,\hat{\varepsilon}^{i}\hat{\varepsilon}_{k}+\sinh(\psi)\underline{K}\,\hat{\varepsilon}_{j}\epsilon^{ji}{}_{k},\quad\underline{K}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (B.13)
represents a twisted boost-rotation operator. Its parameter satisfies $\psi\in(-\infty,\infty)$.
Last, in the case $E=0$, the equations (B.3) decouple, which directly yields the solution
$A^{i}{}^{*}=A^{i},\quad L^{i}{}^{*}=L^{i}+\epsilon^{i}{}_{jk}\varepsilon^{j}A^{k}.$ (B.14)
(2) The basis expansion (A.7)–(A.8) can be used again, where $\Theta^{i}$ is expressed in terms of $M^{i}$ in the case $E\neq 0$ and $A^{i}$ in the case $E=0$. Under the flow (B.4), (B.11), (B.12), (B.14), the resulting expansion then yields the transformations (5.8b) on $r^{i}{}^{*}$ and $v^{i}{}^{*}$.
(3) The flow equation for $t^{*}$ is determined entirely by $\tau^{j}$ which is the same as for the flow arising from the LRL direction vector. Recall that this differential equation can be directly integrated as a quadrature. Then substituting the transformation $r^{i}{}^{*}$ leads to the transformation (5.8a).
Acknowledgements
S.C.A. is supported by an NSERC Discovery grant.
D. Holm and C. Stoica are thanked for helpful remarks.